# Electron-beam annealing of Josephson junctions for frequency tuning of quantum processors

### v1

@@ Abstract
Superconducting qubits are a promising route to achieving large-scale quantum computers. A key challenge in realising large-scale superconducting quantum processors involves mitigating frequency collisions. In this paper, we present an approach to tuning fixed-frequency qubits with the use of an electron beam to locally anneal the Josephson junction. We demonstrate the ability to both increase and decrease the junction barrier resistance. The technique shows an improvement in wafer scale frequency targetting by assessing the frequency collisions in our qubit architecture. Coherence measurements are also done to evaluate the performance before and after tuning. The tuning process utilises a standard electron beam lithography system, ensuring reproducibility and implementation by any group capable of fabricating these Josephson junctions. This technique has the potential to significantly improve the performance of large-scale quantum computing systems, thereby paving the way for the future of quantum computing.
@@ I Introduction
The variability in arises from the Al/AlO/Al stack in the JJs, that are commonly fabricated using double-angled shadow evaporation. Current state-of-the-art shadow-evaporated junctions exhibit a resistance spread between, = 2 % - 3 % at wafer scale (2 and 3 inch) depending on the JJ dimension [9, 10, 14, 15]. Both lateral dimensions of the JJ and the tunneling barrier itself impacts this resistance spread. Dimensional non-uniformities include line edge roughness, and resist stack inhomogeneity [16], resist residues [17], uneven distribution of metal during deposition [14] and variation in critical feature size inherent to tolerance of the lithography tool itself. The tunnelling barrier dimensions and stoichiometry are influenced by the oxidation conditions [18, 19], non-uniform oxide growth across the junction, and several others [20, 21, 22]. To further scale quantum processors, individual qubit frequency targetting must be improved to ensure high gate fidelities, the requirements for this frequency targetting exceeds what is possible with current state-of-the-art as-fabricated Josephson junctions [5, 23].
@@
Electron beam lithography systems widely available in nanofabrication facilities allow this technique to be easily intergrated into fabrication workflows. The technique shows minimal yield reduction, and minimal impact to the qubit coherence metrics. In addition, we demonstrate the ability to tune in a negative direction. We perform wafer level tuning on OQC’s Lucy-generation 8QPU and achieve a resistance spread, 1.5 %. Further, we assess the QPU viability by calculating the frequency-induced collisions before and after e-beam tuning. Lastly, we present the results of qubit coherence measurements showing that the qubits retain their high-coherence after electron-beam annealing.
@@ II Resistance tuning mechanism and procedure using e-beam
The barrier resistance of a JJ has been shown to increase over time, in a process commonly referred to as ‘ageing’. This is an important consideration when fabricating a QPU with precise qubit frequency targets, as those frequencies may drift. The process of JJ ageing appears related to oxygen diffusion or structural alterations in the AlO layer that can be accelerated under elevated temperatures [27]. We leverage this ageing process in order to address the variability in JJ resistances post fabrication.
@@
Figure 1 illustrates the step-by-step procedure to tune the JJ resistance. 1) The process begins with the fabrication of either JJ test structures or wafer-scale QPU’s with lumped elements (resonators and capacitors) fabricated using photolithography, followed by JJs fabricated using EBL and a double angled shadow evaporation technique [34]. Figure 1a shows OQC’s double-sided coaxmon design of the resonator, qubit and JJ elements. 2) We then conduct room temperature resistance measurements on the JJ tunnel barrier to evaluate their resistance distribution , target and yield (Figure 1b and c). Additionally for QPUs, we also derive qubit parameters, such as assessing any potential frequency collisions and calculating the instances where they may occur within the QPUs. To demonstrate a reduction in resistance spread with this technique, we select JJs in the histogram with resistances below the median for tuning, aiming to increase their resistance and narrow the overall distribution (highlighted by the red dashed line in Figure 1c.) 3) These specific JJs are identified and mapped within the sample design layout (Figure 1d) and were assigned the same dose such that the JJs shift as an ensemble to narrow the spread. A more effective approach would be to assign appropriate e-beam doses based on their respective positions in the distribution. 4) Figure 1e shows the tuning being carried out in an EBL system, where the designated JJs are exposed. As the qubits used here are not grounded we use a charge mitigation layer in this processing step to avoid charge build up and any associated yield-reducing processes. 5) After tuning, the resistance measurements are performed again to evaluate the tuned resistance distribution, , where R corresponds to the e-beam annealed resistance, and assess any change in frequency collision metrics. Based on the results, the QPU’s can either be used in a quantum computer, or will undergo another iteration of tuning to further narrow .
@@ III Resistance tuning with applied dose and proximity
IV Reverse resistance tuning
In addition to our primary goal of demonstrating an increase in JJ resistance, we have also been able to achieve JJ resistance reduction. This phenomenon has been observed in studies involving annealing at high temperatures. However, these techniques are non-local and result in a wafer-scale resistance shift [36, 37, 24]. It is suggested that the elevated temperature can induce a phase transition within the JJ barrier oxide from amorphous to polycrystalline, resulting in a reduction in resistance as polycrystalline AlO is shown to be more conductive than its amorphous phase [38, 39]. Alternatively, irradiation with an e-beam rearrangement of thermodynamically less stable bonds in the amorphous matrix prodcuing the polycrystalline phase [40, 41]. The reduction in resistance is evident when a higher e-beam current (5x the current used in resistance increase) is irradiated on the JJ.
Figure 3a shows the resistance distribution (R2) after irradiating a set of JJs with an elevated e-beam current, showcasing a decrease in JJ resistances of approximately 400 when compared to the original median distribution (R1). In this experiment different doses were used however we observe that all the JJs shift together as an ensemble suggesting that the heating profile and proximity effects differ from those typically needed for resistance increase. Figure 3b plots the distribution after a period of ageing (transition from R2 to R3), suggesting that the JJs that have been tuned to a lower resistance are not insensitive to ageing in contrast to previous works [35]. Subsequently, R3 is then tuned to narrow the distribution by increasing the resistance of the JJs below the median (as described in section II). Figure 3c shows that we were able to tune the spread from = 2.35 % to = 1.69 %. This confirms that our technique allows for both resistance reduction and increase in the same JJs. By reducing the resistance, we can increase the accruracy of our global frequency targetting by moving all the JJs down in resistance, and still have the ability to individually tune JJs by increasing the resistance of selected JJs. This allows for the accurate targetting of qubit frequencies followed by precise detuning to reduce collisions. While we have achieved resistance reduction, the precise calibration of resistance and improving the localization of annealing is a current subject of ongoing investigation. Nevertheless, this approach holds promise as it provides a wide range of adjustment possibilities in both directions.
V Wafer level spread reduction
@@
Figure 5a illustrates one of the 8-qubit device selected from Figure 5b, underscoring the initial collision types based on its qubit frequencies before the tuning process. This particular device exhibited a reduction in collisions, from 5 to 0 after the tuning procedure. Figures 5d and 5e provides an evaluation of the 8 qubits on the various collision metrics before and after the tuning process respectively. In the first columns of both figures, the qubit frequencies for the ground state (solid line) and the excited state (dashed line) are presented to assess the C1 criteria. The C1 criteria evaluates whether the frequency gap between neighbouring qubits falls within the straddling regime. The second columns of Figure 5d and 5e provide a corresponding assessment of potential collisions in every qubit based on each criteria. Symbols falling within the red bar signify an impending frequency collision. As expected we observe that the number of S1 collisions (indicated by the blue bar) increase after the tuning process.
VI Coherence measurements after e-beam annealing
@@
VII Conclusion
In conclusion, we have demonstrated the ability to locally tune superconducting qubits via application of an electron-beam directly onto, or in the near-vicinity of, the qubit’s JJ element. Electron-beam tuning was used to locally tune the Josephson junction resistance (correspondingly tuning its frequency). The sensitivity of the junction resistance shift is characterised with respect to both e-beam current and the proximity to the junction, allowing fine control of the junction resistance to meet specific frequency targets. We demonstrate a reduction in resistance (frequency) spread of our qubits by locally tuning Josephson junctions that are lower in resistance than that targetted. We have tuned QPUs at wafer scale, containing a total of 552 Josephson junctions. We reduced the resistance spread from an as-fabricated 3.11 % to 1.48 %, following three e-beam tuning processes. By assessing the expected frequency collisions in our processors we demonstrate that e-beam tuning can be used as a tuning tool to better allocate frequencies in order to reduce frequency collisions and improve expected QPU performance as well as increase the yield of viable QPUs. Finally, we presented results of qubit coherence times and establish that our qubits remain high-coherence after they have undergone the e-beam tuning treatment. With further refinement of this technique, including the optimisation of the negative tuning technique we expect to further the range of tuning that is possible using this process, thus further reducing Josephson junction spread allowing improved qubit frequency allocation. This technique can be performed using tools such as an EBL or a SEM, which allow spatial control of the e-beam. Since these tools are readily available in fabrication facilities, this technique is suitable for integrating into Josephson junction fabrication workflows and processes.

### v2

@@ Abstract
Superconducting qubits are a promising route to achieving large-scale quantum computers. A key challenge in realising large-scale superconducting quantum processors involves mitigating frequency collisions. In this paper, we present an approach to tuning fixed-frequency qubits with the use of an electron beam to locally anneal the Josephson junction. The technique shows an improvement in wafer scale frequency targetting by assessing the frequency collisions in our qubit architecture. Coherence measurements are also done to evaluate the performance before and after tuning. The tuning process utilises a standard electron beam lithography system, ensuring reproducibility and implementation by any group capable of fabricating these Josephson junctions. This technique has the potential to significantly improve the performance of large-scale quantum computing systems, thereby paving the way for the future of quantum computing.
@@ I Introduction
The variability in arises from the Al/AlOx/Al stack in the JJs, that are commonly fabricated using double-angled shadow evaporation. Current state-of-the-art shadow-evaporated junctions exhibit a resistance spread between, = 2 % - 3 % at wafer scale (2 and 3 inch) depending on the JJ dimension [9, 10, 14, 15]. Both lateral dimensions of the JJ and the tunneling barrier itself impacts this resistance spread. Dimensional non-uniformities include line edge roughness, and resist stack inhomogeneity [16], resist residues [17], uneven distribution of metal during deposition [14] and variation in critical feature size inherent to tolerance of the lithography tool itself. The tunnelling barrier dimensions and stoichiometry are influenced by the oxidation conditions [18, 19], non-uniform oxide growth across the junction, and several others [20, 21, 22]. To further scale quantum processors, individual qubit frequency targetting must be improved to ensure high gate fidelities, the requirements for this frequency targetting exceeds what is possible with current state-of-the-art as-fabricated Josephson junctions [5, 23].
@@
Electron beam lithography systems widely available in nanofabrication facilities allow this technique to be easily intergrated into fabrication workflows. The technique shows minimal yield reduction, and minimal impact to the qubit coherence metrics. We perform wafer level tuning on OQC’s Lucy-generation 8QPU and achieve a resistance spread, 1.5 %. Further, we assess the QPU viability by calculating the frequency-induced collisions before and after e-beam tuning. Lastly, we present the results of qubit coherence measurements showing that the qubits retain their high-coherence after electron-beam annealing.
@@ II Resistance tuning mechanism and procedure using e-beam
The barrier resistance of a JJ has been shown to increase over time, in a process commonly referred to as ‘ageing’. This is an important consideration when fabricating a QPU with precise qubit frequency targets, as those frequencies may drift. The process of JJ ageing appears related to oxygen diffusion or structural alterations in the AlOx layer that can be accelerated under elevated temperatures [27]. We leverage this ageing process in order to address the variability in JJ resistances post fabrication.
@@
Figure 1 illustrates the step-by-step procedure to tune the JJ resistance. 1) The process begins with the fabrication of either JJ test structures or wafer-scale QPU’s with lumped elements (resonators and capacitors) fabricated using photolithography, followed by JJs fabricated using EBL and a double angled shadow evaporation technique [34]. Figure 1a shows OQC’s double-sided coaxmon design of the resonator, qubit and JJ elements. 2) We then conduct room temperature resistance measurements on the JJ tunnel barrier to evaluate their resistance distribution , target and yield (Figure 1b and c). Additionally for QPUs, we also derive qubit parameters, such as assessing any potential frequency collisions and calculating the instances where they may occur within the QPUs. To demonstrate a reduction in resistance spread with this technique, we select JJs in the histogram with resistances below the median for tuning, aiming to increase their resistance and narrow the overall distribution (highlighted by the red dashed line in Figure 1c.) 3) These specific JJs are identified and mapped within the sample design layout (Figure 1d) and were assigned the same dose such that the JJs shift as an ensemble to narrow the spread. A more effective approach would be to assign appropriate e-beam doses based on their respective positions in the distribution. 4) Figure 1e shows the tuning being carried out in an EBL system, where the designated JJs are exposed. As the qubits used here are not grounded we use a charge mitigation layer in this processing step to avoid charge build up and any associated yield-reducing processes. 5) After tuning, the resistance measurements are performed again to evaluate the tuned resistance distribution, , where Re corresponds to the e-beam annealed resistance, and assess any change in frequency collision metrics. Based on the results, the QPU’s can either be used in a quantum computer, or will undergo another iteration of tuning to further narrow .
@@ III Resistance tuning with applied dose and proximity
IV Wafer level spread reduction
@@
Figure 4a illustrates one of the 8-qubit device selected from Figure 4b, underscoring the initial collision types based on its qubit frequencies before the tuning process. This particular device exhibited a reduction in collisions, from 5 to 0 after the tuning procedure. Figures 4d and 4e provides an evaluation of the 8 qubits on the various collision metrics before and after the tuning process respectively. In the first columns of both figures, the qubit frequencies for the ground state (solid line) and the excited state (dashed line) are presented to assess the C1 criteria. The C1 criteria evaluates whether the frequency gap between neighbouring qubits falls within the straddling regime. The second columns of Figure 4d and 4e provide a corresponding assessment of potential collisions in every qubit based on each criteria. Symbols falling within the red bar signify an impending frequency collision. As expected we observe that the number of S1 collisions (indicated by the blue bar) increase after the tuning process.
V Coherence measurements after e-beam annealing
@@
VI Conclusion
In conclusion, we have demonstrated the ability to locally tune superconducting qubits via application of an electron-beam directly onto, or in the near-vicinity of, the qubit’s JJ element. Electron-beam tuning was used to locally tune the Josephson junction resistance (correspondingly tuning its frequency). The sensitivity of the junction resistance shift is characterised with respect to both e-beam current and the proximity to the junction, allowing fine control of the junction resistance to meet specific frequency targets. We demonstrate a reduction in resistance (frequency) spread of our qubits by locally tuning Josephson junctions that are lower in resistance than that targetted. We have tuned QPUs at wafer scale, containing a total of 552 Josephson junctions. We reduced the resistance spread from an as-fabricated 3.11 % to 1.48 %, following three e-beam tuning processes. By assessing the expected frequency collisions in our processors we demonstrate that e-beam tuning can be used as a tuning tool to better allocate frequencies in order to reduce frequency collisions and improve expected QPU performance as well as increase the yield of viable QPUs. Finally, we presented results of qubit coherence times and establish that our qubits remain high-coherence after they have undergone the e-beam tuning treatment. With further refinement of this technique we expect to further the range of tuning that is possible using this process, thus further reducing Josephson junction spread allowing improved qubit frequency allocation. This technique can be performed using tools such as an EBL or a SEM, which allow spatial control of the e-beam. Since these tools are readily available in fabrication facilities, this technique is suitable for integrating into Josephson junction fabrication workflows and processes.